OpenStudy (anonymous):

i chose c am i correct? 1. Which function is an odd function? (Points : 1) f(x) = 0.8x^3 f(x) = 5x^2 f(x) = 0.7x^2 + 3 f(x) = 3x - 1

OpenStudy (anonymous):

@hartnn what do you think?

hartnn (hartnn):

for an odd function f(-x) = -f(x)

hartnn (hartnn):

so, just replace 'x' by -x in your functions, what do u get ?

OpenStudy (anonymous):

f(x) = 0.8(-x)^3 f(-x) = (5-x)^2 f(-x) = 0.7(-x)^2 + 3 f(=x) = 3(-x) - 1

hartnn (hartnn):

so whats (-x)^3= ... ?

OpenStudy (anonymous):

-x *-x * -x = (+)x^3

hartnn (hartnn):

2 negatives become positive but 3 negatives is negative only

OpenStudy (anonymous):

but the exponent wouldn't it be -x multiplied by itself 3 times?

hartnn (hartnn):

yes

hartnn (hartnn):

-x *-x * -x = (-)x^3

OpenStudy (anonymous):

but its separated like this?: (-x * -x) *-x

hartnn (hartnn):

x^2*-x = -x^3

OpenStudy (anonymous):

ok so , then my answer is b or c

hartnn (hartnn):

no, you need the - for a, f(-x)=-f(x) so its odd

OpenStudy (anonymous):

its odd if (-x) is not equal to f(x) or -f(x) right?

OpenStudy (anonymous):

i have them mixed up

hartnn (hartnn):

its odd if it is =-f(x)

OpenStudy (anonymous):

so a = 0.8 * (-x) ^3 = -0.8 depending on what x is?

hartnn (hartnn):

what!?

OpenStudy (anonymous):

for it to be odd it has to = -f(x)

hartnn (hartnn):

f(x) = 0.8x^3 is odd, because f(-x) = -f(x)

OpenStudy (anonymous):

oh, so i really just had to focus on f(x) like the choices of a and b because i was only given f(-x) not in slope intercept form